OpenStudy (anonymous):

I will fan and medal!! Bulk food at a grocery store is sold by the pound. The more you buy, the more you pay. But the price per pound is constant. Use this graph to determine the price per pound of bulk granola. A graph titled, Bulk Granola is shown. The x-axis is labeled Weight (pounds) the y-axis is labeled Cost (in U S dollars). Points are plotted at two comma five, four comma ten, and six comma fifteen. A line is drawn through those points. A. fifty cents per pound B. two dollars per pound C. two dollars and fifty cents per pound D. three dollars and fifty cent

OpenStudy (anonymous):

@Vocaloid

Vocaloid (vocaloid):

price per pound = slope

OpenStudy (anonymous):

it's A?

OpenStudy (anonymous):

no C?

OpenStudy (anonymous):

so y is dollars, x is pounds.

OpenStudy (anonymous):

All you need is slope ^^^^

OpenStudy (texaschic101):

slope formula : slope = (y2 - y1) / (x2 - x1) (2,5) and (4,10) slope = (10 - 5) / (4 - 2) can you finish and find the slope ?

OpenStudy (anonymous):

no :(

OpenStudy (anonymous):

2.5

OpenStudy (texaschic101):

slope = (10 - 5) / (4 - 2) slope = 5/2....turned to decimal = 2.50 per lb

OpenStudy (anonymous):

C

OpenStudy (texaschic101):

yes :)

OpenStudy (anonymous):

there you go!